OpenStudy (goformit100):

Find all integers n(positive, negative or zero) for which n^2 + n + 41 is a perfect square.

OpenStudy (experimentx):

write it up as n^2 + n + 41 = m^2 n + 41 = (m-n)(m+n) note that m-n>0 and m-n < 7 put m-n = 1,2, ...,6 and try to solve for for integer value of n. if you get integer value of n, then that's your answer for n. finally you should get this result http://www.wolframalpha.com/input/?i=n%5E2+%2B+n+%2B+41++%3D+m%5E2+Interger+solution

OpenStudy (goformit100):

Thank you sir